I need to know how to solve for x when the question is like 14,583=5.77x-8.4
I would really appriciate your help. I am in desperet need.
thanks

Answer by pridgea(6)

14583+8.4=5.77x- 8.4+8.4; 14591.4/5.77 =5.77x/5.77; 2528.84=x